Evaluate the following residues. A. (10 points) Res ( rez B. (10 points) Res (2.44 +3:1)
Answer #1

4) Red ( 127) f(z) - el poles of fcz] =) 24_16=0 (23 4 2 - 2 - 4 (2²+4)=0 z = 12 Z = 2i Resid z = +2,-2,21, -2i Such pole of2 +3 P 2+3 212-1² poles of fcz) are z=0, 2-1 at zoo is a pole D order) at 2:1 is a pole of orders ResCfid) - cm 243 2(22) - (